LOW TEMPERATURE CHEMICAL REACTOR

\underline{MARIN PICHLER}, WILLIAM C. STWALLEY, Department of Physics, University of Connecticut, Storrs CT 06269.

A new approach for production of very cold molecules is described. This includes an experimental system for producing and observing Cs^{∗} + H_{2} → CsH + H chemical reactions below 10 K. The ground state Cs atoms are trapped and cooled down to few hundred microKelvin in a standard magneto-optical trap (MOT), and then excited to the 7P_{1/2} or 7P_{3/2} state. A cold para H_{2} beam (up to 10 K) collides with Cs^{∗} atoms. Due to the reaction energetics, the CsH molecules produced by the 7P_{1/2} reaction must be in the lowest rovibrational state. Detection techniques to be employed include LIF and REMPI.